dahgirl2:

Use the function f(x) = 2x2 − x − 10 to answer the questions. Part A: Completely factor f(x). Part B: What are the x-intercepts of the graph of f(x)? Show your work. Part C: Describe the end behavior of the graph of f(x). Explain. Part D: What are the steps you would use to graph f(x)? Justify that you can use the answers obtained in Part B and Part C to draw the graph.

BestBoiKirby:

@dude

dude:

Hey! Do you know how to factor?

dahgirl2:

um no nr

dude:

That's okay! We want to start out by finding which numbers multiplied will all up to 10 Could you list them?

dahgirl2:

1*10 2*5 5*2

Aqual:

@Salmon

dude:

Yeah! Now we have to find which pair adds up to -1. Do you know how to do that?

dahgirl2:

no how do i do that

dude:

How would you add up to -1 from each pair? Try: -1 + 10 = ? 1 + -10 = ? -5+2 = ? -2+5 = ?

dahgirl2:

9 -9 -3 3

dude:

Oh hold on, I did not notice the 2 coefficient The 2 coefficient changes the equation Try inserting 2 into each would-be factor to see if it makes the equation (Usually if the number is smaller, add to where the smaller factor will be multiplied by this coefficient) Ie. -5+2 = -3 (-5 is smaller you'd preferrably want it to multiply with this) And such, it seems to be true \((2x-5)(x+2) = 2x^2+4x-5x-10=> 2x^2-x-10\)

dahgirl2:

ok thank u